Trying to solve this equation. Have four choices on worksheet but cant get the right answer. the problem is (7x-7)squared=36. The choices are 13/7,0
29/7 -13/7,-1/7 13/7,1/7 thanks

the next problem is 2n squared = -8n-1 thank you again

Answer by kapilsinghi(68)

Trying to solve this equation. Have four choices on worksheet but cant get the right answer. the problem is (7x-7)squared=36. The choices are 13/7,0
29/7 -13/7,-1/7 13/7,1/7 thanks
(7x-7)^2 = 36
taking squareroot on both sides
7x-7=6 or 7x-7=-6
7x=6+7 or 7x=-6+7
7x=13 or 7x=1
x=13/7 or x=1/7

kapilsinghi123@gmail.com